the wieghts in pounds of a set of newborn triplets are 4 3/4 pounds,4 5/8 pounds,and 4 7/16 punds.order the weights from least to greatest
Answer by mananth(16949)

the wieghts in pounds of a set of newborn triplets are 4 3/4 pounds,4 5/8 pounds,and 4 7/16 punds.order the weights from least to greatest
4 3/4 = 19/4
4 5/8 = 37/8
4 7/16 = 71 /16
The idea is to convert the denominator in each fraction to 16 so that all the fractions have the same denominator.
The value of the fraction does not change if you divide and multiply the fraction by the same number
Once the fractions have the same denominator the fractions can be arranged from least to highest by noting the numerator value.
19/4 * 4/4 = 76/16
37/8 * 2/2 = 74/ 16
71/16 = 71/16

4 3/4 = 19/4-----------------76/16 greatest
4 5/8 = 37/8 ----------------74/16
4 7/16 = 71 /16--------------71/16 least